With the increase in price of tomato from Rs. 50 per kg to Rs. 100 per kg in blue market the demand for tomato has gone down from 40 kg to 30 kg. Calculate the demand elasticity.
1
Expert's answer
2018-06-27T08:47:08-0400
The price increases from $50 to $100.
Therefore % change = (100-50)/50 = 1 (100%)
1 = 100% (1 *100)
Quantity fell by (30-40)/40 = – 0.25 (25%)
Therefore PED = (% change in Q.D.)/(% change in Price)=(-25)/100
Therefore PED = -0,25
If the elasticity is less than unity, inelastic demand